A Treasury bond with the longest maturity (30 years) has an ask price quoted at 102.9375. The coupon rate is 3.90 percent, paid semiannually. What is the yield to maturity of this bond? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places.)
Answer #1

Let Par Value =100
Price =102.9375
Number of Periods =30*2 =60
Semi annual coupon =3.90%*100/2 =1.950
YTM using excel function =2*RATE(60,19.50,-102.9375,100) =3.74%